OpenStudy (anonymous):

Given this vectorial subspace in R4 generated by these vectors: V=< (0, 0, 0, 1), (0, 0, 1, 2), (1, 2, 3, 4)> Provide the base for it

OpenStudy (anonymous):

@FSU

OpenStudy (anonymous):

tbh im not sure

OpenStudy (anonymous):

@blake007

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

ow :(

OpenStudy (anonymous):

@saifoo.khan please help?

OpenStudy (amistre64):

base, or basis?

OpenStudy (anonymous):

basis i believe

OpenStudy (anonymous):

sorry I speak spanish so I get troubles with translating

OpenStudy (amistre64):

a basis is a minimum set of linearly independant column vectors that provides a way to approach all the vectors in the the subspace. does that make sense as a definition?

OpenStudy (anonymous):

yes, I believe I need three vectors that are a system generator i think and also linearly independent?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

I'm just not sure how to approach he problem in order to obtain those three vectors I was trying to do just elimination in a matrix but i couldnt

OpenStudy (amistre64):

echelon comes to mind but im not quite sure if we are given row vectors or column vectors at the moment. quite frankly, i recall doing something like seting up the matrix A = v1, v2, v3 and getting it into row echelon form

OpenStudy (anonymous):

I got a another group of vectors and i used them as rows didnt know they had to specify

OpenStudy (anonymous):

it does not specify if its row or column , can i use it as i wish?

OpenStudy (amistre64):

thats the part i cant recall at the moment :) last one a worked on said something to the effect: v1, v2, v3 form a basis of a subspace if for some abc consants av1 + bv2 + cv3 = 0 is not the trivial solution, a=b=c=0

OpenStudy (anonymous):

so thats a (0,0,0,1) + b (0,0,1,2) + c(1,2,3,4) =0 ?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

let me review for a moment to make sure my memory isnt failing me on this point

OpenStudy (anonymous):

sure, no problem

OpenStudy (amistre64):

av1 + bv2 + cv3 = 0 is ONLY the trivial case. for example, in R^3 we have the basis v1 = 001 v2 = 010 v3 = 100 the only way a linear combination of these vectors add up to 0 is if a=b=c=0

OpenStudy (anonymous):

this is R^4

OpenStudy (amistre64):

it applies to any vector space, now my question is, are we spose to create the basis using the vectors in V only, or are we to create it a basis that can obtian at least the vectors that are in V

OpenStudy (amistre64):

http://www.math.odu.edu/~bogacki/cgi-bin/lat.cgi well ill be, theres a webpage i just found that walks you thru it

OpenStudy (anonymous):

is asking only to find a basis or the vectorial space

OpenStudy (anonymous):

oh wow thanks ! :D

OpenStudy (amistre64):

it cleared up my concerns thats for sure. i knew your set V was linearly independant, i echeloned it, and was of the impression that the vectors themselves formed the basis they wanted.

OpenStudy (anonymous):

I cannot use those vectors as the basis you mean?

OpenStudy (amistre64):

i thought that was linking to the page i worked out, its this one one the ment it turns out that v1,v2,v3 form the basis itself

OpenStudy (anonymous):

thank you.

OpenStudy (amistre64):

youre welcome